# {textblock*}

### v1

@@ 1 Introduction
It has been known for a long time that dS solutions are hard to realize in string theory. More recently, this has been turned into a Swampland Conjecture against stringy dS models [1, 2]. Renewed scrutiny of the leading stringy dS models [3, 4] has then led to the preliminary conclusion that serious control issues exist [5, 6, 7, 8, 9, 10, 11, 12]. See also [13, 14, 15, 16, 17, 18] for recent work aimed at improving these leading dS constructions. Independently of such stringy issues, arguments against the consistency of dS space have been discussed for decades, see e.g. [19, 20, 21, 22, 23, 24, 25, 26].
Let us for the moment assume that string theory or even quantum gravity in general do indeed abhor dS space. If so, the question is which particular feature of this cosmological solution makes it problematic. A fairly obvious possibility is the presence of cosmological horizons [27, 28, 29]. The option of forbidding them has recently been revived in the Swampland context [30, 31]. We will call the corresponding condition the ‘Horizon Criterion’, to be specified more precisely momentarily. The point of the present paper is to focus on this condition and to investigate what it implies for the allowed geometries and scalar potentials. Of course, this is closely connected with the swampland-type question about accelerated expansion in asymptotic regimes of string compactifications [32, 33, 34] and, more specifically, about precise bounds on the underlying asymptotic potentials [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. For earlier studies of eternal cosmic acceleration in the string theory context see e.g. [47, 48, 49, 50, 51, 52, 53]. However, as we discuss at the end of Sect. 2, the corresponding ‘no asymptotic acceleration conjecture’ is hard to make suitably precise or phrase as a fundamental quantum gravity constraint. In our subsequent study of the ‘Horizon Criterion’ we will be particularly interested in spacetimes with boundaries, including but not limited to models with localized gravity [54, 55, 56, 57], as recently discussed in the Swampland context in [58]. [^1] The idea of dynamical spacetime boundaries or ‘End-of-the-World’ branes first appeared in [59] and are expected to be ubiquitous in quantum gravity thanks to the Cobordism conjectures [60]. For a selection of recent related studies see [61, 62, 63, 64, 65, 66].
@@
Cosmological horizons:
Consider a spacetime $\mathcal{M}$ and an inertial observer with worldline $\sigma$ reaching infinite proper time. If there exist points $p\in\mathcal{M}$ such that $\sigma$ does not intersect the domain of future influence222 We refrain from writing ‘forward light cone’ since we do not want black hole horizons to fall under our definition. In the black hole case, we assume that due to unitary evaporation the region of future influence of points inside the horizon is not limited by the singularity. of $p$ we will say that ‘$p$ is screened from the observer by a cosmological horizon.’ This horizon is hence the boundary between points which are screened and those which are not. Here, an observer is considered inertial if its worldline is a geodesic. In case the observer is localized to the spacetime boundary, the worldline must be a geodesic with respect to the metric induced on the boundary.
Based on this definition, one may try to formulate a ‘Horizon Criterion’ by claiming that string theory does not allow for solutions with cosmological horizons. To illustrate this, consider $d$-dimensional FLRW spacetimes,
@@
This implies that light-rays traverse asymptotically a finite coordinate range with respect to $ds_{d-1}^{2}$, establishing the existence of a cosmological horizon. A necessary condition for horizon formation is that spacetime undergoes eternal accelerated expansion, characterized by $\ddot{a}(\tau)>0$ for $\tau>\tau_{0}$ 333 Nevertheless, as will be argued in Sect. 2, we expect that horizons and not eternal acceleration represent the fundamentally problematic feature of de Sitter-like solutions.
If the energy density of an FLRW universe is dominated by a scalar field $\phi_{\rm FLRW}$ with potential $V(\phi_{\rm FLRW})\sim\exp({-\gamma\phi_{\rm FLRW}})$, then asymptotic cosmic acceleration requires [36]
@@
which in this case also leads to the presence of a horizon (as emphasized in the present context in [30, 31]). Thus, flat potentials are ruled out. As we will see, generalizing this logic to spacetimes with boundaries is non-trivial and, since boundaries are in general part of string theory solutions, more care is required in formulating a horizon criterion.
@@
The study of swampland conjectures in the presence of boundaries is interesting both as a matter of principle and because boundaries are expected to be ubiquitous in string theory [60]. If the bulk is AdS and the ETW-brane tension is appropriately adjusted, then gravity localizes at this so-called ‘UV brane’ [55] and one is effectively dealing with a lower-dimensional Minkowski-space theory. This continues to hold for ‘de-tuned’ brane-tension, with the lower-dimensional effective theory being correspondingly AdS or Minkowski [56, 57]. One may view this as a cutoff-version [67, 68, 69, 70, 71] of AdS/CFT [72]: The $d$-dimensional AdS bounded by a UV brane corresponds to a $(d\!-\!1)$-dimensional gravitational theory coupled to a CFT with UV-cutoff. The latter represents the bulk degrees of freedom.
@@
Following this logic, one concludes that the swampland constraints on such spacetimes are just the ordinary swampland constraints of a $(d\!-\!1)$-dimensional gravity + CFT system. Nevertheless, it is of course interesting to check this explicitly, as proposed recently in [58]. Specifically in our context, one may consider a model with ETW-brane-localized scalar field $\phi$. Its potential is then subject to the constraints of the dS-conjecture or, if we assume horizons to be the problematic feature of dS spaces, to the Horizon Criterion.
@@
An interesting observation can be made by considering a flat ETW brane, i.e. a UV-brane with tension tuned as in the original Randall-Sundrum (RS) model [55]. We note that this model is not in the swampland since an explicit string-theoretic realization exists [67]. The Penrose diagram of this model is shown in Fig. 1. More precisely, the RS model directly corresponds to the Poincaré patch of AdS, colored gray in the figure, together with the brane shown in red. But this spacetime is only part of global AdS.
@@
Figure 1: The Penrose diagram of $d$-dimensional AdS with a flat Randall-Sundrum UV brane embedded (red line). The gray region corresponds to a portion of a Poincaré patch of AdS.
@@
The UV-brane of the RS model intersects the asymptotic boundary at infinite Poincaré time, which is however at finite global time. The region outside the Poincaré patch is inaccessible for an observer living on the brane. Thus, according to the above definition there exists a horizon. In other words, the RS model realizes a spacetime with cosmological horizon while not being in the swampland. A swampland conjecture excluding horizons can hence not be upheld. If one wants to maintain that horizons are the underlying reason for why de Sitter space is problematic in string theory, a refined version of the horizon criterion is required.
The brane-brane horizon criterion:
@@
For universes with boundary, one can distinguish between horizons for observers living in the bulk or on the brane. Consider first an inertial bulk observer having a cosmological horizon. We will call the intersection of this horizon with the bulk a bulk-bulk horizon and its intersection with the ETW brane a bulk-brane horizon. Similarly, for an observer on the ETW brane who sees a horizon, we will call the intersection of the horizon with the bulk a brane-bulk horizon and its intersection with the ETW brane a brane-brane horizon.
@@
According to these definitions, the RS spacetime only has a brane-bulk horizon whereas the dS branes of [56, 57] realize a spacetime with brane-bulk and brane-brane horizons. A refined statement of the horizon criterion would then be that consistent theories should not allow for solutions which possess brane-brane horizons.
@@
The condition for the scalar potential of brane-localized scalar that arises from the horizon criterion is already known for ETW branes bounding AdS spacetimes: One simply makes use of AdS/CFT with cutoff and applies (3). We will hence focus on studying the situation when ETW branes bound flat spacetimes. Although gravity does not localize on such branes, they can still provide an expanding lower-dimensional geometry potentially possessing brane-brane horizons. [^4] One might naively think that decays to nothing realize an ETW boundary with horizon. However, such decays can only occur for non-SUSY spacetimes. As one expects all Minkowski spaces in the string landscape to be supersymmetric, the only interesting case is the decay of AdS. But unstable AdS decays completely within one AdS-time, such that every observer dies in finite time and the concept of a horizon does not make sense. We focus on ETW branes whose energy density is set by a brane-localized scalar $\phi$ and analyze conditions on the potential $V(\phi)$ that guarantee or exclude brane-brane horizons. The dynamics of ETW branes of this type bounding AdS space has recently been considered in [73] in the context of AdS/CFT correspondence.
@@
$V(\phi)=\begin{cases}\frac{c^{2}}{\phi^{2}}\,,\qquad\text{with}\qquad c<c_{\rm crit}=\frac{1}{\sqrt{2(4d-9)}}&\phi\text{ living on the brane}\\ V_{0}e^{-\lambda\phi}\,,\qquad\text{with}\qquad\lambda<\lambda_{\rm crit}=\frac{1}{\sqrt{d-2}}&\phi\text{ living in the bulk}\end{cases}$ (4)
@@
Furthermore, employing an autonomous system analysis (see [74] for a review on autonomous systems in cosmology), we study the ‘critical cases’, corresponding to potentials that to leading order take the form (4) with $c=c_{\rm crit}\,,\lambda=\lambda_{\rm crit}$ and derive conditions on corrections to the potential that guarantee the existence of brane-brane horizons. We also study the corresponding result for ordinary FLRW universes, refining the result (3).
This paper is structured as follows. In Sect. 2, we discuss the brane-brane horizon criterion in more detail. In Sect. 3, we consider a scalar field $\phi$ only living on the ETW brane and derive a general condition for brane-brane horizons to exist, in analogy to (2). We subsequently study which scalar field potentials realize them. In Sect. 4, we consider $\phi$ to be a modulus in the bulk. As before, we derive a condition analogous to (2) and study its implications on the scalar field potential on the boundary. In App. C, we study solutions that marginally do (not) realize horizons in more detail.
2 Preliminary Reflections on the Horizon Criterion
@@
One may be concerned that inertial observers on an ETW brane are very different from inertial bulk observers and that a ‘Horizon Criterion’ should only apply to the latter. As a defense against such criticism, we note that there exist situations in which the UV resolution of an ETW brane is a boundary-free higher-dimensional spacetime, such that the inertial ETW brane observer is in fact identical to an inertial bulk observer. The most prominent example of this is Witten’s bubble of nothing [75]. It describes the decay of the Kaluza-Klein vacuum $\mathbb{M}^{1,3}\times S^{1}$, i.e. of 4d Minkowski space times an $S^{1}$, to nothing. [^5] See [76] for a string-theoretic construction of a bubble of nothing, establishing their existence in UV complete quantum gravity, and [63] for a proposal for how bubbles of nothing could naturally be part of the 4d type-IIB flux landscape. In the Lorentzian geometry, the $S^{1}$ shrinks to vanishing size over an $S^{2}\subset\mathbb{R}^{3}\subset\mathbb{R}^{1,3}$. This $S^{2}$ and hence the ball of ‘nothing’ inside it grow as a function of time. From the 4d perspective, one has an expanding ETW brane with negative tension, located on this $S^{2}$ [63]. Interestingly, the worldline of an inertial observer on the ETW brane, i.e. on the (5d smooth) locus where the $S^{1}$ reaches zero size, is also a geodesic in the 5d metric. Such an an observer is then both an inertial ETW-brane and an inertial bulk observer. In such situations, the more standard Horizon-Criterion for 5d observers implies the horizon criterion for 4d brane observers.
@@
As a further concern, one may note that Rindler observers exist even in Minkowski space. They have a horizon and, as result, forbidding geometries in which certain observers have horizons appears unmotivated. Indeed, consider for example a detuned RS model in which the UV brane has a tension above the critical value. This brane accelerates outward, inducing both brane-brane and brane-bulk horizons. Similarly, a Rindler observer accelerates and induces a brane-bulk horizon. However, there is a key difference: The acceleration of the brane-bound observer originates in the ETW brane tension. This is fundamentally different a Rindler observer, who needs to carry the energy for the acceleration with them, such that eternal acceleration appears impossible, even in principle. We hence do not view the existence of (idealized) Rindler observers as a strong argument against our proposal.
Potential counterexamples from decays of AdS or Minkowski space:
One might naively think that horizons can be trivially obtained by considering the decay of an AdS spacetime to lower-energy AdS or to nothing. Such a proposal [77] has been developed to explain the accelerated expansion we currently observe in our universe. The domain wall or ETW brane appearing in this process may then induce brane-brane horizons. However, such models assume unstable AdS, and it is well-known that the latter decays completely in finite time (of order the AdS-time) [78]. In particular, infinitely many bubbles form in the infinite-volume region near the AdS-boundary. Their collisions prevent the worldline of a brane-localized observer to reach the AdS-boundary, such that no Randall-Sundrum-type horizon can form.

### v2

# and the Horizon Criterion

@@ 1 Introduction
It has been known for a long time that dS solutions are hard to realize in string theory. More recently, this has been turned into a Swampland Conjecture against stringy dS models Danielsson:2018ztv, Obied:2018sgi. Renewed scrutiny of the leading stringy dS models Kachru:2003aw, Balasubramanian:2005zx has then led to the preliminary conclusion that serious control issues exist Carta:2019rhx, Gao:2020xqh, Junghans:2022exo, Gao:2022fdi, Junghans:2022kxg, Hebecker:2022zme, Schreyer:2022len, Schreyer:2024pml. See also Demirtas:2021nlu, Hebecker:2020ejb, Krippendorf:2023idy, McAllister:2024lnt, Lanza:2024uis, Chauhan:2025rdj for recent work aimed at improving these leading dS constructions. Independently of such stringy issues, arguments against the consistency of dS space have been discussed for decades, see e.g. Ford:1984hs, Mottola:1984ar, Antoniadis:1985pj, Tsamis:1992sx, Goheer:2002vf, Polyakov:2007mm, Dvali:2014gua, Dvali:2017eba.
Let us for the moment assume that string theory or even quantum gravity in general do indeed abhor dS space. If so, the question is which particular feature of this cosmological solution makes it problematic. A fairly obvious possibility is the presence of cosmological horizons Fischler:2001yj, Hellerman:2001yi, Bedroya:2019snp. The option of forbidding them has recently been revived in the Swampland context Hebecker:2023qke, Andriot:2023wvg. We will call the corresponding condition the ‘Horizon Criterion’, to be specified more precisely momentarily. The point of the present paper is to focus on this condition and to investigate what it implies for the allowed geometries and scalar potentials. Of course, this is closely connected with the swampland-type question about accelerated expansion in asymptotic regimes of string compactifications Ooguri:2018wrx, Hebecker:2018vxz, Andriot:2018mav and, more specifically, about precise bounds on the underlying asymptotic potentials Rudelius:2021oaz, Rudelius:2021azq, Rudelius:2022gbz, Andriot:2022xjh, Calderon-Infante:2022nxb, Marconnet:2022fmx, Shiu:2023nph, Shiu:2023fhb, Cremonini:2023suw, Freigang:2023ogu, VanRiet:2023cca, Marconnet:2025vhj (see also Garg:2018reu). For earlier studies of eternal cosmic acceleration in the string theory context see e.g. Townsend:2003fx, Ohta:2003pu, Chen:2003dca, Andersson:2006du, Tsujikawa:2013fta, Russo:2018akp, Grimm:2019ixq. However, as we discuss at the end of Sect. 2, the corresponding ‘no asymptotic acceleration conjecture’ is hard to make suitably precise or phrase as a fundamental quantum gravity constraint. In our subsequent study of the ‘Horizon Criterion’ we will be particularly interested in spacetimes with boundaries, including but not limited to models with localized gravity Randall:1999ee, Randall:1999vf, Kaloper:1999sm, Karch:2000ct, as recently discussed in the Swampland context in Anastasi:2025puv. [^1] The idea of dynamical spacetime boundaries or ‘End-of-the-World’ branes first appeared in Jourjine:1983du and are expected to be ubiquitous in quantum gravity thanks to the Cobordism conjectures McNamara:2019rup. For a selection of recent related studies see Buratti:2021yia, Blumenhagen:2022mqw, Friedrich:2023tid, Sugimoto:2023oul, Angius:2024pqk, Muntz:2024joq, Barbon:2025vvh.
@@
Cosmological horizons: Consider a spacetime $\mathcal{M}$ and an inertial observer with worldline $\sigma$ reaching infinite proper time. If there exist points $p\in\mathcal{M}$ such that $\sigma$ does not intersect the domain of future influence [^2] We refrain from writing ‘forward light cone’ since we do not want black hole horizons to fall under our definition. In the black hole case, we assume that due to unitary evaporation the region of future influence of points inside the horizon is not limited by the singularity. of $p$ we will say that ‘$p$ is screened from the observer by a cosmological horizon.’ This horizon is hence the boundary between points which are screened and those which are not. Here, an observer is considered inertial if its worldline is a geodesic. In case the observer is localized to the spacetime boundary, the worldline must be a geodesic with respect to the metric induced on the boundary.
Based on this definition, one may formulate the ‘Horizon Criterion’ by claiming that string theory does not allow for solutions with cosmological horizons. To illustrate this, consider $d$-dimensional FLRW spacetimes,
@@
This implies that light-rays traverse asymptotically a finite coordinate range with respect to $ds_{d-1}^{2}$, establishing the existence of a cosmological horizon. A necessary condition for horizon formation is that spacetime undergoes eternal accelerated expansion, characterized by $\ddot{a}(\tau)>0$ for $\tau>\tau_{0}$ .  [^3] Nevertheless, as will be argued in Sect. 2, we expect that horizons and not eternal acceleration represent the fundamentally problematic feature of de Sitter-like solutions.
If the energy density of an FLRW universe is dominated by a scalar field $\phi_{\rm FLRW}$ with potential $V(\phi_{\rm FLRW})\sim\exp({-\gamma\phi_{\rm FLRW}})$, then asymptotic cosmic acceleration requires Rudelius:2021azq
@@
which in this case also leads to the presence of a horizon (as emphasized in the present context in Hebecker:2023qke, Andriot:2023wvg). Thus, flat potentials are ruled out. As we will see, generalizing this logic to spacetimes with boundaries is non-trivial and, since boundaries are in general part of string theory solutions, more care is required in formulating a ‘Refined Horizon Criterion’.
@@
The study of swampland conjectures in the presence of boundaries is interesting both as a matter of principle and because boundaries are expected to be ubiquitous in string theory McNamara:2019rup. If the bulk is AdS and the ETW-brane tension is appropriately adjusted, then gravity localizes at this so-called ‘UV brane’ Randall:1999vf and one is effectively dealing with a lower-dimensional Minkowski-space theory. This continues to hold for ‘de-tuned’ brane-tension, with the lower-dimensional effective theory being correspondingly AdS or Minkowski Kaloper:1999sm, Karch:2000ct. One may view this as a cutoff-version Verlinde:1999fy, Gubser:1999vj, Giddings:2000mu, Duff:2000mt, Arkani-Hamed:2000ijo of AdS/CFT Maldacena:1997re: The $d$-dimensional AdS bounded by a UV brane corresponds to a (d1)(d\!-\!1)-dimensional gravitational theory coupled to a CFT with UV-cutoff. The latter represents the bulk degrees of freedom.
@@
Following this logic, one concludes that the swampland constraints on such spacetimes are just the ordinary swampland constraints of a (d1)(d\!-\!1)-dimensional gravity + CFT system. Nevertheless, it is of course interesting to check this explicitly, as proposed recently in Anastasi:2025puv. Specifically in our context, one may consider a model with ETW-brane-localized scalar field $\phi$. Its potential is then subject to the constraints of the dS-conjecture or, if we assume horizons to be the problematic feature of dS spaces, to the Horizon Criterion.
@@
An interesting observation can be made by considering a flat ETW brane, i.e. a UV-brane with tension tuned as in the original Randall-Sundrum (RS) model Randall:1999vf. We note that this model is not in the swampland since an explicit string-theoretic realization exists Verlinde:1999fy. The Penrose diagram of this model is shown in Fig. 1. More precisely, the RS model directly corresponds to the Poincaré patch of AdS, colored gray in the figure, together with the brane shown in red. But this spacetime is only part of global AdS.
@@
Figure 1: The Penrose diagram of $d$-dimensional AdS with a flat Randall-Sundrum UV brane embedded (red-green dashed line). The gray region corresponds to a portion of a Poincaré patch of AdS. The worldline $\sigma$ of a brane-bound observer is illustrated in green. It coincides with the location of the RS brane in this illustration, which is why the spacetime boundary is illustrated as a red-green dashed line. One observes that there exists a point p whose future domain of influence does not intersect $\sigma$.
@@
The UV-brane of the RS model intersects the asymptotic boundary at infinite Poincaré time, which is however at finite global time. The region outside the Poincaré patch is inaccessible for an observer living on the brane. Thus, according to the above definition there exists a horizon. In other words, the RS model realizes a spacetime with cosmological horizon while not being in the swampland. A swampland conjecture excluding horizons can hence not be upheld. [^4] A horizon also exists in the UV completion. To see this, recall that the Verlinde model Verlinde:1999fy may be thought of as a compactification on $T^{6}/\mathbb{Z}_{2}$ with a stack of 16 D3 branes sitting at one point of the compact space. Replacing this stack by a warped throat, $\mathbb{R}_{+}\times S^{5}$, one discovers that the near-brane geometry is that of an $AdS_{5}\times S^{5}$ compactification from 10d to 5d. This $AdS_{5}\times S^{5}$ region may be viewed as the bulk of an RS model. The rest of the $T^{6}/\mathbb{Z}_{2}$ plays the role of the UV brane. Now, observers living in this compact space (e.g. bound to one of the O3-planes) see a horizon precisely in the sense of Fig. 1. Namely, signals from the part of the $AdS_{5}\times S^{5}$ geometry above the gray region can not reach them. If one wants to maintain that horizons are the underlying reason for why de Sitter space is problematic in string theory, a refined version of the Horizon Criterion is required.
The Refined Horizon Criterion:
@@
For universes with boundary, one can distinguish between horizons for observers living in the bulk or on the brane.
Brane and bulk horizons: For a bulk observer with cosmological horizon, we will call the intersection of this horizon with the bulk a bulk-bulk horizon. Its intersection with the ETW brane will be called its bulk-brane horizon. Similarly, for an observer on the ETW brane who sees a horizon, we will call the intersection of the horizon with the bulk the brane-bulk horizon and its intersection with the ETW brane the brane-brane horizon.
@@
According to these definitions, the RS spacetime only has a brane-bulk horizon whereas the dS branes of Kaloper:1999sm, Karch:2000ct realize a spacetime with brane-bulk and brane-brane horizons. We can now formulate a refined version of the Horizon Criterion:
Refined Horizon Criterion: Consistent theories of quantum gravity do not allow for spacetime geometries exhibiting bulk-bulk or brane-brane horizons.
@@
The condition for the scalar potential of brane-localized scalar that arises from this Refined Horizon Criterion is already known for ETW branes bounding AdS spacetimes: One simply makes use of AdS/CFT with cutoff and applies (3). We will hence focus on studying the situation when ETW branes bound flat spacetimes. Although gravity does not localize on such branes, they can still provide an expanding lower-dimensional geometry potentially possessing brane-brane horizons. [^5] One might naively think that decays to nothing realize an ETW boundary with horizon. However, such decays can only occur for non-SUSY spacetimes. As one expects all Minkowski spaces in the string landscape to be supersymmetric, the only interesting case is the decay of AdS. But unstable AdS decays completely within one AdS-time, such that every observer dies in finite time and the concept of a horizon does not make sense. We focus on ETW branes whose energy density is set by a brane-localized scalar $\phi$ and analyze conditions on the potential $V(\phi)$ that guarantee or exclude brane-brane horizons. The dynamics of ETW branes of this type bounding AdS space has recently been considered in Fujiki:2025yyf in the context of AdS/CFT correspondence.
@@
[…]
@@
Furthermore, employing an autonomous system analysis (see Bahamonde:2017ize for a review on autonomous systems in cosmology), we study the ‘critical cases’, corresponding to potentials that to leading order take the form (4) with $c=c_{\rm crit}\,,\lambda=\lambda_{\rm crit}$ and derive conditions on corrections to the potential that guarantee the existence of brane-brane horizons. We also study the corresponding result for ordinary FLRW universes, refining the result (3).
This paper is structured as follows. In Sect. 2, we discuss the Refined Horizon Criterion in more detail. In Sect. 3, we consider a scalar field $\phi$ only living on the ETW brane and derive a general condition for brane-brane horizons to exist, in analogy to (2). We subsequently study which scalar field potentials realize them. In Sect. 4, we consider $\phi$ to be a modulus in the bulk. As before, we derive a condition analogous to (2) and study its implications on the scalar field potential on the boundary. In App. C, we study solutions that marginally do (not) realize horizons in more detail.
2 Preliminary Reflections on Horizon Criteria
In this section, we critically reflect on Horizon Criteria as discussed in the Introduction.
2.1 On different types of observers
In this subsection, we discuss the notion of an observer in more detail: We show that bulk observers in the UV theory may sometimes become brane observers from an IR perspective and critically discuss the special role of Rindler observers.
@@ 2.1 On different types of observers
One may be concerned that inertial observers on an ETW brane are very different from inertial bulk observers and that a horizon criterion should only apply to the latter. As a defense against such criticism, we note that there exist situations in which the UV resolution of an ETW brane is a boundary-free higher-dimensional spacetime, such that the inertial ETW brane observer is in fact identical to an inertial bulk observer. The most prominent example of this is Witten’s bubble of nothing Witten:1981gj. It describes the decay of the Kaluza-Klein vacuum $\mathbb{M}^{1,3}\times S^{1}$, i.e. of 4d Minkowski space times an $S^{1}$, to nothing. [^6] See GarciaEtxebarria:2020xsr for a string-theoretic construction of a bubble of nothing, establishing their existence in UV complete quantum gravity, and Friedrich:2023tid for a proposal for how bubbles of nothing could naturally be part of the 4d type-IIB flux landscape. In the Lorentzian geometry, the $S^{1}$ shrinks to vanishing size over an $S^{2}\subset\mathbb{R}^{3}\subset\mathbb{R}^{1,3}$. This $S^{2}$ and hence the ball of ‘nothing’ inside it grow as a function of time. From the 4d perspective, one has an expanding ETW brane with negative tension, located on this $S^{2}$ Friedrich:2023tid. Interestingly, the worldline of an inertial observer on the ETW brane, i.e. on the (5d smooth) locus where the $S^{1}$ reaches zero size, is also a geodesic in the 5d metric. Such an observer is then both an inertial ETW-brane and an inertial bulk observer. In such situations, the standard Horizon Criterion for 5d observers implies the Refined Horizon Criterion as stated in the Introduction.
@@
As a further concern, one may note that Rindler observers exist even in Minkowski space. They have a horizon and, as result, forbidding geometries in which certain observers have horizons appears unmotivated. Indeed, consider for example a detuned RS model in which the UV brane has a tension above the critical value. This brane accelerates outward, inducing both brane-brane and brane-bulk horizons. Similarly, a Rindler observer accelerates and induces a brane-bulk horizon. However, there is a key difference: The acceleration of the brane-bound observer originates in the ETW brane tension. By contrast, the acceleration of a Rindler observer requires an energy source, which the observer must carry with them and which will eventually be exhausted. As a result, eternal acceleration of the Rindler observer seems impossible – even in principle. We hence do not view the existence of (idealized) Rindler observers as a strong argument against our proposal.
2.2 Potential counterexamples
In this subsection, we address potential counterexamples to the proposed Refined Horizon Criterion.
One might naively think that horizons can be trivially obtained by considering the decay of an AdS spacetime to lower-energy AdS or to nothing. Such a proposal Banerjee:2018qey has been developed to explain the accelerated expansion we currently observe in our universe. The domain wall or ETW brane appearing in this process may then induce brane-brane horizons. However, such models assume unstable AdS, and it is well-known that the latter decays completely in finite time (of order the AdS-time) Horowitz:2007pr. In particular, infinitely many bubbles form in the infinite-volume region near the AdS-boundary. Their collisions prevent the worldline of a brane-localized observer to reach the AdS-boundary, such that no Randall-Sundrum-type horizon can form.